OpenStudy (ray_sihota):

If i have a payment of 805.02 and I have made ten payments , how do i find out what I still owe?

OpenStudy (ray_sihota):

is it a formula , or tvm? or is it the Amortization Function

OpenStudy (triciaal):

not enough information? you have made a total of 805.02 * 10 = 8050.20 how much did you owe to start with? example the amount of a loan how long was the "contract" how many payments did you agree to make?

OpenStudy (ray_sihota):

Judith bought a condo for 115000. She paid 40,000 down and agreed to make equal payments at the end of every month for25 year. The rate was 13.25% compounded monthly I worked out the payments to 860.02 (sorry about that) B) after 10 years of payments , how much will judith owe.

OpenStudy (ray_sihota):

I converted 10 year into 120 months

OpenStudy (triciaal):

loan amount / amount financed = 75000 time left = 25 - 10 = 15 years interest rate = 13.25 compounded monthly =(13.25/12) Compound interest = Pe^rt P = principal r = interest rate as % and t = time in years value = principal + interest Alternate you might need to do the present value of an annunity sorry, It's been a while since I did finance

OpenStudy (ray_sihota):

that's what I was thinking , thank you anyways for your assistance.